OpenStudy (anonymous):

What is the value of the expression –9 – 7 + 5? A. 3 B. –7 C. 11 D. –11

OpenStudy (anonymous):

THE � ARE THESE -

Nnesha (nnesha):

what is - 9 - 7 = ???

OpenStudy (anonymous):

-16

Nnesha (nnesha):

ok so what -16 + 5 = ???

OpenStudy (anonymous):

-11

Nnesha (nnesha):

yes :)

OpenStudy (anonymous):

:) thanks

Nnesha (nnesha):

my pleasure :)

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

wait can u help me with some more

Nnesha (nnesha):

sure

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

What is the value of the expression |x + y| + |z| when x = –6, y = 4, and z = –13?

OpenStudy (anonymous):

the �� are -

OpenStudy (anonymous):

i think the answer is -15

OpenStudy (anonymous):

i have to type the anwser on this one

OpenStudy (anonymous):

answer*

Nnesha (nnesha):

suppose to be positive 15